OpenStudy (anonymous):

WILL FAN AND MEDAL The incomplete work of a student to solve an equation is shown below: Step 1: 2x + 8 = 2 Step 2: ? Step 3: x = −6 ÷ 2 Step 4: x = −3 What is the missing Step 2? A. 2x=-6 B. 2x=6 C. 2x=10 D. 2x=-10 B.

OpenStudy (plasmataco):

A.

OpenStudy (plasmataco):

2x=2-8

OpenStudy (solomonzelman):

do we report those that give answers in a manner that you did?

OpenStudy (plasmataco):

you subtract both sides by 8.

OpenStudy (plasmataco):

?

OpenStudy (solomonzelman):

(I didn't report, sateline is that is why you came ))

OpenStudy (plasmataco):

I don't really understand.

OpenStudy (anonymous):

Yeah, thanks for the answer but you aren't supposed to give direct answers. But again thank you.

OpenStudy (plasmataco):

oh. sorry.

OpenStudy (ybarrap):

2x + 8 = 2 What do you need to do to get rid of that 8 ? Then what would result when you do that? 2x + 8 \(\color{red}{- 8}\) = 2 \(\color{red}{- 8}\)

OpenStudy (anonymous):

It would be 2x=-6?

OpenStudy (anonymous):

Thank you for explaining the step. :)

OpenStudy (ybarrap):

You're welcome